OpenStudy (anonymous):

In the equation x – 4y = 2, for what value of x will y = 0?

OpenStudy (whpalmer4):

\[x-4y=2\]You want to find the value of \(x\) when \(y=0\), so put \(y=0\) in the equation and solve for \(x\): \[x-4(0) = 2\]\[x-0=2\]\[x=\]

OpenStudy (anonymous):

so its -2 right?

OpenStudy (whpalmer4):

\[x-0=2\]From that you get \(x=-2\)?

OpenStudy (whpalmer4):

Does the equation work if you substitute in \(x=-2,y=0\)? Let's find out: \[x-4y=2\]\[-2-4(0) = 2\]\[-2-0=2\]\[-2=2\]Uh, no.

OpenStudy (anonymous):

so its 2 0r -2

OpenStudy (whpalmer4):

You have some number, \(x\). Here's an equation: \[x - 0 = 2\] Translated into words, that reads: "take your number, \(x\), and subtract 0 from it. The result is 2." What is the only value that \(x\) can have to make that sentence true?

OpenStudy (anonymous):

2-0=2

OpenStudy (anonymous):

right?

OpenStudy (whpalmer4):

Yes. \[x-4y=2\]\[2-4(0)=2\]\[2-0=2\]\[2=2\]